OpenStudy (anonymous):

Auditions for a play in New York were held with 120 men and women trying out for the lead singer position. If twice the number of men less three times the number of women is equal to 20, how many women auditioned? Need help setting it up

OpenStudy (anonymous):

m + w = 120 3w - 2m = 20

OpenStudy (anonymous):

maybe i'm doing it wrong but i got 52... and its not one of the options

OpenStudy (anonymous):

@telltoamit

OpenStudy (anonymous):

m =68 , w = 52 thats what i am getting

OpenStudy (anonymous):

me too but look 44 40 76 80

OpenStudy (anonymous):

umm, wel , the language of the question isnt very clear i think they hv taken 2m - 3w = 20

OpenStudy (anonymous):

so now idk should i just guess?